OpenStudy (anonymous):

how to do equation of circle centered at 5,2 and a radius of 2?

OpenStudy (solomonzelman):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (solomonzelman):

\[(x-5)^2+(y-2)^2=4\]

OpenStudy (anonymous):

so what do you do with the x and y?

OpenStudy (solomonzelman):

you don;t do anything with x and y, you just plug in (h,k) inside the formula.

OpenStudy (solomonzelman):

\[\huge\color{blue}{(x-h)^2+(y-k)^2=r^2}\] (h,k) is the center of the circle.

OpenStudy (anonymous):

ok so the x and y are cancelled out?

OpenStudy (solomonzelman):

no, they remain the same, but h and k change into 2 and 5 and r^2 becomes 4 because you are substituting the coordinate of the center of the circle and the radius that you are given into the formula. MAKES SENSE?

OpenStudy (anonymous):

yea, thanks. also one quick one also.

OpenStudy (anonymous):

what is one similarity between a cylinder and rectangular prism?

OpenStudy (solomonzelman):

 Well really, nothing, other than that they are shapes with a volume or 3D shapes, and both of their volumes would be written in cubic units.

OpenStudy (anonymous):

Ok thanks. That was a question my teacher asked me and i couldnt figure it out.

OpenStudy (solomonzelman):

Anytime!